### v1

# Activation Steering in Neural Theorem Provers

Abstract
Large Language Models (LLMs) have shown promise in proving formal theorems using proof assistants like Lean. However, current state of the art language models struggle to predict next step in proofs leading practitioners to use different sampling techniques to improve LLMs capabilities. We observe that the LLM is capable of predicting the correct tactic; however, it faces challenges in ranking it appropriately within the set of candidate tactics, affecting the overall selection process. To overcome this hurdle we use activation steering to guide LLMs responses to improve the generations at the time of inference. Our results suggest that activation steering offers a promising lightweight alternative to specialized fine-tuning for enhancing theorem proving capabilities in LLMs, particularly valuable in resource-constrained environments.
1 Introduction
Interactive proof assistants such as Lean de Moura et al. (2015), Isabelle Wenzel et al. (2008), and Coq Barras et al. (1999) enable the formal verification of mathematical proofs and software by leveraging specialized programming languages Avigad (2023); Ringer et al. (2019). Neural theorem proving, which integrates neural language models with interactive proof assistants, has emerged as a promising approach to automating formal reasoning First et al. (2023); Polu & Sutskever (2020b); Polu et al. (2022); Yang et al. (2023b); Welleck (2023). This integration is mutually beneficial: proof assistants enforce formal correctness, while language models assist in proof construction by predicting and suggesting logical steps. A central challenge in this setting is tactic prediction—determining the appropriate next step at each proof state.
In this work, we investigate activation steering Panickssery et al. (2024); Turner et al. (2024); Lucchetti & Guha (2024) as a technique to enhance tactic prediction in Llemma Azerbayev et al. (2024a) and InternLM2 Ying et al. (2024a). These language models are designed for theorem proving by training and fine-tuning on mathematical data. Activation steering is an inference-time model editing method that modifies a model’s internal representations to guide its behavior toward desired outputs. We propose its application in refining tactic selection, aiming to improve both the accuracy and interpretability of proof automation. By systematically influencing LLMs’ reasoning process, our approach enables structured interventions that enhance model-driven theorem proving, leading to more reliable and controllable predictions.
We present an approach for steering tactic selection from a pair of prompts $(p_{1},p_{2})$ that contain a LEAN state $s$ to generate the next step (or tactic) $t$. However, the LLM successfully predicts $t$ for $p_{1}$ but mispredicts for $p_{2}$. In each pair $p_{2}$ is natural data and $p_{1}$ is synthetically generated using $p_{2}$. We systematically add the attributes and a high level of structure the proof should follow. These additional attributes guide the model to a specific and more grounded chain of thought to follow while predicting the tactics. These abstractions over the proof help the LLM to do critical decision making while predicting the next step.
2 Related Work
2.1 Formal Theorem Proving
Formal theorem proving encodes theorems and proofs in a machine-verifiable format, ensuring correctness through rigid logical rules. A key component of this field is Interactive Theorem Proving (ITP), where humans collaborate with proof assistants such as Isabelle Wenzel et al. (2008), Lean de Moura et al. (2015), and Coq Barras et al. (1999) to formally verify proofs. These assistants allow users to express theorems in higher-order logic and construct verifiable proofs.
In Lean de Moura et al. (2015), proofs are built using tactics, which either solve a goal or decompose it into sub-goals.
2.2 Proofstep Generation with Large Language Models
Generating intermediate proof steps is a fundamental challenge in theorem proving, particularly in tactic-based automated theorem provers (ATPs). Early neural approaches (Whalen, 2016; Huang et al., 2019; Bansal et al., 2019; Paliwal et al., 2020; Sanchez-Stern et al., 2020) framed proofstep generation as a classification task, employing models like TreeLSTM and RNN to predict tactics and their arguments. ASTactic (Yang & Deng, 2019) later introduced a grammar-constrained decoder for structured tactic generation.
Recent advances leverage large language models (LLMs) for proof generation, casting tactic prediction as an extitauto-regressive sequence modeling problem. GPT-$f$ (Polu & Sutskever, 2020a) pioneered this approach by training transformers only with decoders to generate structured proof steps. Baldur (First et al., 2023) extended this by producing entire proofs, while POETRY (Wang et al., 2024) adopted a recursive decomposition strategy. Other works (Szegedy et al., 2021; Tworkowski et al., 2022; Welleck et al., 2022; Jiang et al., 2022; Yang et al., 2023a) integrate the selection of premises with tactic prediction, employing retrieval-augmented methods and constrained decoding to improve the coherence of the proof.
Lean-Star Lin et al. (2024) introduced Self-Taught Reasoning, incorporating Chain-of-Thought Wei et al. (2023) reasoning before each tactic to generate synthetic data for fine-tuning LLMs via self-play Chen et al. (2024). Our work leverages these randomly sampled data points from Lean-STaR-base as natural data.
2.3 Mechanistic Interpretability
Previous work has focused on localizing and editing factual associations within transformers (Meng et al., 2022) and probing hidden representations for high-level knowledge (Li et al., 2024b; Dong et al., 2023). Such studies perform implicit evaluations of model ability, complementing explicit benchmarks (Dong et al., 2023). A key technique in mechanistic interpretability is activation patching (Vig et al., 2020; Variengien & Winsor, 2023), which modifies model activations to influence outputs. This research has suggested the existence of task vectors (Hendel et al., 2023; Ilharco et al., 2022)—representations encoding abstract task information. Activation steering has been employed to mitigate model deceitfulness and sycophancy (Rimsky et al., 2023; Li et al., 2024b), further supporting the presence of task vectors. Steering is based on the linear representation hypothesis (Park et al., 2023), which posits that concepts exist as directions in the embedding space of the model.
For theorem proving, mechanistic interpretability provides insights into how LLMs represent logical structures and reasoning processes. By dissecting these representations, we can identify failure cases, refine tactic prediction, and enhance proof generation. We hypothesize that effective steering transforms activations to align the model’s reasoning trajectory with a more structured and verifiable direction.
2.4 Model Steering
Activation-based interventions can directly influence the language model output during inference Dathathri et al. (2019); Subramani et al. (2022). Recent studies demonstrate that activation steering enhances truthfulness , mitigates sycophancy, and improves instruction-following Stolfo et al. (2024), as well as type prediction in code Lucchetti & Guha (2024).
Building on these ideas, we investigate steering vectors in the context of theorem proving. Following prior work Burns et al. (2024); Turner et al. (2024); Arditi et al. (2024); van der Weij et al. (2024), we compute steering vectors based on input pairs differing by a specific feature—here, the presence or absence of synthetic metadata. Unlike previous studies that focused on broad linguistic properties such as sentiment and style, our approach seeks to refine the logical inference pathways of LLMs.
Figure 1: Steering Vectors are computed as difference of activations of $p_{1}$ and $p_{2}$
For theorem proving, activation steering provides a mechanism to guide the model towards structured reasoning, improving its ability to generate valid proof steps. By leveraging task vectors, we aim to shift model activations to align with correct logical deductions, thereby enhancing both proof coherence and model interpretability. This intervention is particularly valuable in interactive theorem proving, where fine-grained control over reasoning steps can lead to more reliable and verifiable proofs.
3 Methodology
3.1 Constructing Steering Dataset
Model Choice  We build steering datasets for 7B parameter Llemma Azerbayev et al. (2024b) and 7B parameter InternLM2 Ying et al. (2024b). These models are trained to generate formal theorems and code, which is important for the tactic prediction task.
Source Dataset We constructed steering pairs from a randomly sampled subset $S$ from Lean-STaR data. The subset consists of approximately ten thousand unique proof stages and tactics. We treat this subset as natural prompt data $p_{2}$ and generate a new set $p_{1}$ by adding reasoning steps in $p_{2}$ representing a lean stage $l$. This generates a set of pairs ($p_{1}$, $p_{2}$) which we then prompt InternLM2 to predict the next tactic $t_{1}$ and $t_{2}$ respectively. We then take pairs where $t_{1}\neq t_{2}$ creating a subset $s$ of prompt pairs. We then validate tactics $t_{1},t_{2}\in s$ with Lean Prover to generate $s^{\prime}$.
This process generates a quadruple $\{p_{1},p_{2},t_{1},t_{2}\}$ where $t_{1}$ and $t_{2}$ are valid tactics for a lean stage $l$. We assume that $t_{1}$ is a more optimal tactic for $l$.
Generating $p_{1}$ from $p_{2}$ In Fig. 2, we illustrate the process of improving theorem-proving LLMs using StepBackReasoning Zheng et al. (2024). Initially, the model is given a prompt asking it to predict the next tactic in a Lean 4 proof. Without additional reasoning guidance, the model produces an incorrect output, such as $rfl$. which does not align with the required proof strategy. To address this, we introduce a step-back reasoning mechanism. First, we extract the proof state from the given Lean prompt, as shown in the StepBackAbstraction module. This abstraction step helps identify the core mathematical principles underlying the theorem. Using GPT-4, we generate step-back reasoning prompts, which explicitly highlight key mathematical structures, such as exponentiation in a division monoid and properties like associativity and commutativity. These enriched insights are then incorporated into the original theorem-proving prompt, providing the model with a more structured understanding of the problem. As a result, when the theorem-proving LLM is queried again, it produces the correct output $rwpow_{m}ul$ which correctly applies the relevant exponentiation rule. This process demonstrates how structured reasoning about the problem enhances logical coherence and guides LLMs toward more accurate proof generation in Lean 4.
Figure 2: Steering Vectors are computed as difference of activations of $p_{1}$ and $p_{2}$
3.2 Constructing Steering Vectors
Given dataset of steering pairs and tactics $\{p_{1},p_{2},t_{1},t_{2}\}\in s^{\prime}$ and a model $M$, we apply a forward pass to every $M(p_{1})$,$M(p_{2})$ to collect values of the residual stream vector on queries at the last token of the input layer $\ell\in\{1,...,L\}$. We isolate the internal representation corresponding to the reasoning by computing the difference in residual stream vectors $v_{1,\ell}$ and $v_{2,\ell}$. More formally, we compute the vector $u_{\ell}$ representing the direction of the steering at layer $\ell$:
$\mathbf{u}_{\ell}=\frac{\mathbf{v}_{\ell}}{\|\mathbf{v}_{\ell}\|},\quad\text{where}\quad\mathbf{v}_{\ell}=\frac{1}{N}\sum_{i}^{N}\left(\mathbf{p}_{1,\ell,i}-\mathbf{p}_{2,\ell,i}\right)$
Averaging our different proof states to capture activation values most closely associated with the structured reasoning step independent of the query. The calculation of the direction of the steering is carried out using the representations in the last token of the input, which effectively encapsulates the behavior of the model not only for the next token prediction task, but also for the entire generation following Todd et al. (2024); Scalena et al. (2024). After identifying steering direction, we compute steering vector by re-scailing unit vector $u_{\ell}$ by a coefficient $c$. We use a systematic scaling approach where the value of c is selected to ensure that residual stream activations are assigned to their mean value on inputs that contain the structured reasoning steps. In particular, we compute a new example with residual stream values $p^{\prime}$ at a given token.
$c=\bar{z}-p_{\ell}^{\prime}u_{\ell},\quad\text{where}\quad\bar{z}=\frac{1}{N}\sum_{i}^{N}p_{1,i,\ell}u_{\ell}.$
The steering vector $cu_{\ell}$ is then added to the corresponding residual stream layer and the forward pass is resumed with the updated residual stream value $\tilde{x}_{\ell}^{\prime}=x_{\ell}^{\prime}+cu_{\ell}$. This procedure is carried out at a single layer across all token positions, motivated by previous findings that show models tend to deviate from instructions as they generate more tokens Stolfo et al. (2024); Li et al. (2024a)
4 Results and Analysis
Setup We evaluated the technique using Best First Search. It is one of the most popular methods to evaluate the theorem-proving ability of a language model Polu & Sutskever (2020b); Yang et al. (2023b); Azerbayev et al. (2023); Lin et al. (2024). For a given language model $M$, we keep all unexpanded states $s_{i}$; each time we expand the best state $s_{i}$ and use the language model to sample $S$ net tactics $a_{i,1...S}$ for the current state $s_{i}$. Following standard practice Polu & Sutskever (2020b); Yang et al. (2023b); Welleck & Saha (2023); Lin et al. (2024) we assume the state with maximum negative log-probabilities is the ”best” state. Specifically, we select state $s_{i}$ with maximum $\sum_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),...,(s_{i-1},a_{i-1})$ is proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average log probability of each generated token. We expand upto $N$ states and we get successful proofs search when we reach any proof state with no goals.
Model Decoding N K S MiniF2F
GPT-3.5 (FEW-SHOT) SAMPLING 50 1 1 2.8%
GPT-4 (FEW-SHOT) SAMPLING 50 1 1 11.9%
LLEMMA-7B SEARCH 50 1 32 26.2%
INTERNLM2-7B SEARCH 50 1 32 30.3%
INTERNLM2-7B (SFT) SEARCH 50 1 32 30.7%
LEAN-COT (INTERNLM2-7B) SAMPLING 50 32 1 27.0%
LEAN-COT (INTERNLM2-7B) SEARCH 50 1 32 25.4%
LEAN-STAR (INTERNLM2-7B) SAMPLING 50 32 1 29.1%
LEAN-STAR (INTERNLM2-7B) SEARCH 50 1 32 26.2%
CORPA (WITH GPT-4) CUSTOMIZED - 60 1 29.9%
OURS (LLEMMA-7B) SAMPLING 50 32 1 28.1%
OURS (LLEMMA-7B) SEARCH 50 1 32 26.3%
OURS (INTERNLM2-7B) SAMPLING 50 32 1 32.4%
OURS (INTERNLM2-7B) SEARCH 50 1 32 26.8%
Table 1: Pass rates on the miniF2F-test dataset with Lean. This table shows the pass rates of previous works and our work. S is the number of tactics attempted at each expanded node (assumed to be 1 in sampling), and K is the total number of search or sampling attempts per problem.
Dataset We evaluated our technique on MiniF2F benchmark Zheng et al. (2022). Which consists of 244 theorems in lean 4. We use the same evaluation setting as previous works Yang et al. (2023b); Welleck & Saha (2023); Ying et al. (2024a).
Sampling We evaluate two different decoding strategies: sampling and search. Sampling involves drawing multiple proof steps stochastically based on the model’s output distribution, promoting diversity in the generated proofs. In contrast, search incorporates structured exploration techniques to improve proof discovery. Specifically, we consider two widely used search methods: beam search and best-first search.
Beam Search. Beam search maintains a fixed number $k$ of proof trajectories at each step, selecting the top-ranked candidates based on the model’s confidence scores. By preserving multiple plausible proof paths instead of greedily committing to the highest-confidence step, beam search mitigates early pruning errors and allows exploration of alternative reasoning chains. However, the trade-off between beam width and computational cost remains a key consideration: while larger beams enhance robustness, they also introduce significant overhead.
Best-First Search. Best-first search prioritizes proof states according to a heuristic function, typically based on model confidence or learned value estimates. Unlike depth-first or breadth-first strategies, best-first search expands the most promising proof state first, dynamically adjusting the exploration process. In our experiments, we observe that combining best-first search with sampling yields notable improvements. We hypothesize that this effect arises because traditional reranking, despite boosting the likelihood of high-reward tactics, may suffer from premature convergence to suboptimal proof paths. In contrast, best-first search, when coupled with sampling, allows for imperfect scoring, thereby encouraging broader exploration and improved intermediate state discovery.
4.1 Main Results
Our main results are reported in Table 1. Steering the model’s activation significantly improves performance over the base model. Notably, we observe that steering markedly increases pass rates when using Best First Search with sampling. We hypothesize that this improvement occurs because reranking may be too narrowly focused—potentially getting trapped in local optima—even though it boosts log probabilities for tactics that follow the highest reward path. In contrast, combining sampling with Best First Search allows for imperfect scoring, which in turn enables the exploration of nodes that lead to better intermediate states.
4.2 Ablations
Random Steering Vectors
A potential validity concern with any intervention involving activation patching is that the observed improvements might not stem from genuine performance enhancements but rather from activating fallback mechanisms McGrath et al. (2023); Lucchetti & Guha (2024). For instance, patching could merely introduce noise into the embedding space, inadvertently triggering alternative pathways that lead to the desired outcome. This phenomenon complicates the interpretability of both patches and steering vectors. To examine this, we conduct an experiment using a randomly generated steering vector (denoted as ”Random” in Table 2). Our findings show that even random steering achieves a nonzero accuracy, albeit significantly lower than that of our computed steering vectors. We hypothesize that this residual accuracy arises due to backup circuits. Nevertheless, the substantially higher performance of our computed steering vectors suggests that our approach induces meaningful transformations toward the correct target.
Model Decoding Random Steering
LLEMMA-7B SAMPLING 22.7% 28.1%
SEARCH 19.2% 26.3%
INTERNLM-7B SAMPLING 21.4% 32.4%
SEARCH 18.9% 26.8%
Table 2: Pass rates with randomized vectors and steering vectors.
5 Conclusion
We investigate activation steering for tactic prediction by making language models adhere to structured reasoning approaches in theorem proving. We find that by constructing steering pairs using synthetic metadata and natural proof states, we can construct effective steering vectors that improve tactic selection. Our experiments show that steering vectors enhance model performance beyond random interventions and generalize well across different theorem-proving strategies. The effectiveness of our steering approach demonstrates the existence of underlying reasoning pathways that can be systematically influenced within language models. Activation steering proves to be a powerful technique for improving model performance on formal reasoning tasks where fine-tuning may be impractical or resource-intensive. As language models continue to evolve in their theorem-proving capabilities, activation steering may serve as a lightweight alternative to specialized fine-tuning approaches. Rather than training separate models for different theorem-proving strategies, the same base model could be adapted through steering vectors that guide its reasoning process as needed. This could be particularly valuable for interactive theorem proving environments where computational resources are limited. Our reasoning-based steering vectors, for example, could provide an efficient way to enhance proof assistants, particularly useful for applications like automated tactic suggestion. In future work, we aim to study the underlying mechanisms in language models responsible for structured mathematical reasoning. We further wish to explore how reasoning-focused steering vectors may generalize to open-ended theorem proving tasks and investigate their potential for improving proof search strategies.
6 Future Works
Our work opens several promising research directions for improving theorem proving with activation steering. From a theoretical perspective, we aim to investigate the geometric properties of steering vectors in the context of formal reasoning, studying how different model layers represent logical structures and how steering vectors interact with these representations. Understanding these properties could lead to more efficient steering methods and deeper insights into how LLMs encode mathematical concepts. On the practical side, several extensions could enhance theorem proving systems, including the development of adaptive steering mechanisms that dynamically adjust based on proof state complexity, the investigation of steering vector composition for handling compound mathematical concepts, and the integration of steering with existing proof search heuristics to improve exploration efficiency. Scalability remains a challenge, and future work should explore techniques for reducing the computational overhead of steering during inference, methods for distilling steering vectors while maintaining their effectiveness, and approaches for generalizing steering vectors across different mathematical domains. Our findings suggest that activation steering could become a powerful tool for enhancing LLM-based theorem provers, particularly in resource-constrained environments where fine-tuning is impractical. We believe that exploring these directions will lead to more robust and efficient theorem proving systems.